OpenStudy (anonymous):

Which equations in point-slope form are equations of the line that pass through the points (−7, 5) and (−4, 1)?

OpenStudy (anonymous):

@EmilyD22 @shifuyanli

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you got the slope? it is \[\frac{5-1}{-7-(-4)}\]

OpenStudy (misty1212):

better knows as \(-\frac{4}{3}\)

OpenStudy (misty1212):

then put in point slope form doesn't matter which point you use, could be \[y-5=-\frac{4}{3}(x+7)\] or \[[y-1=-\frac{4}{3}(x+4)\]